OpenStudy (anonymous):

Find the value of the discriminant for the following quadratic equation: 2x2 + 6x + 5 = 0

OpenStudy (anonymous):

Discriminant i.e. D=b^2-4ac

OpenStudy (rock_mit182):

\[\Delta = b ^{2}-4ac\]

OpenStudy (rock_mit182):

Where: \[ax ^{2}+bx+c =0\]

OpenStudy (anonymous):

here a=2 b=6 c=5 therefore D= 6^2-4(2X5) D= 36-40 D=-6

OpenStudy (rock_mit182):

if \[\Delta > 0 = \] real roots